# Potential trace inequalities via a Calderón-type theorem

### v1

@@ Abstract.
We establish an approach to trace inequalities for potential-type operators based on an appropriate modification of an interpolation theorem due to Calderón. We develop a general theoretical tool for establishing boundedness of notoriously difficult operators (such as potentials) on certain specific types of rearrangement-invariant function spaces from analogous properties of operators that are easier to handle (such as fractional maximal operators). The key ingredient for the development of the theory is the initial pair of endpoint estimates for the easier operator whose pivotal example is based on a two-weight inequality of Sawyer. Among various applications we obtain a generalization of the celebrated trace inequality involving the Riesz potential and the Hausdorff content by Korobkov and Kristensen.
@@
Trace inequality, Hausdorff content, interpolation of operators, Calderón theorem, rearrangement-invariant spaces, Lorentz spaces, potentials
@@ 1. Introduction and statement of main results
using Hahn-Banach one obtains that
@@
for every $f\in L^{p,1}(\mathbb{R}^{n})$. One can remove the $\approx$ in this statement about Banach space duality by utilizing a norm on $L^{1}(\mathcal{H}^{n-\alpha p}_{\infty})$ in place of the above quasi-norm, for example with a Choquet integral involving the corresponding dyadic Hausdorff content. Now, while Theorem A asserts that the left-hand side of (1.6) is bounded from above by a constant multiple of $\|f\|_{L^{p,1}(\mathbb{R}^{n})}^{p}$, [Kristensen-Korobkov, Theorem 1.2] asserts the same for the right-hand side of (1.6). Therefore, the formulations are indeed equivalent.
@@
The starting point of this paper is an observation concerning the connection of the result of Korobkov and Kristensen and a classical result of E. Sawyer [Sawyer1, Sawyer2], combined with the same duality principle as in Remark 1.1. To this end, we recall that in his papers on one and two weight estimates, Sawyer proved (see also [AdamsChoquet, pp. 28–29]) the following: For $\beta\in(0,n)$ and $1<q<\frac{n}{\beta}$ there exists a constant $C_{7}=C_{7}(\beta,q,n)>0$ such that
@@
For $\beta=\alpha$ and $q=p$, the inequality (1.4) shows that the estimate for the fractional maximal function in (1.8) is better than that for the Riesz potential of the same order in (1.1), while as discussed in the preceding, there is no hope to control the integral of the Riesz potential with respect to the Hausdorff content of this order.
@@
Thus, for any $\alpha\in(0,n)$ and $1<p<\frac{n}{\alpha}$, one may choose $\beta\in(\alpha,n)$ such that $q=p\frac{\alpha}{\beta}>1$. Then for $f=\chi_{E}$ and $\theta=\alpha/\beta$, (1.9) yields
@@
In this work we develop a framework for this principle, when a bound for a good operator like the fractional maximal operator can easily be translated into a bound for a (comparatively) bad operator like the Riesz potential. Our results are in the general context of rearrangement-invariant function spaces, which preempts the question of bounds for bad operators with that of necessary and sufficient conditions for bounds for good operators on these spaces. The general context of rearrangement-invariant function spaces provides an unifying theory for function spaces such as Lebesgue spaces, Lorentz spaces, or Orlicz spaces, to name a few.
@@
As we will see, this results in significant differences in the theory we develop from the classical one. Several other types of various nonstandard versions of Calderón’s theorem can be found in literature, see e.g. [Bae:22, GP-Indiana:09, Mal:12], but none of the known ones can be used for our purposes.
@@
We say that a quasilinear operator $T$ defined on $(L^{p}+L^{q,\infty})(\mathcal{R},\mu)$ and taking values in $\mathfrak{M}_{0}(\mathcal{S},\nu)$ is $(p,q)$-sawyerable if $T$ satisfies:
@@
Recall that an operator $T$ defined on a linear space $X\subseteq\mathfrak{M}_{0}(R,\mu)$ and taking values in $\mathfrak{M}_{0}(S,\nu)$ is quasilinear if there is a constant $k\geq 1$ such that
@@
for every $f,g\in X$ and every scalar $\alpha$. We say that $T$ is sublinear if it is quasilinear with $k=1$.
@@
    Every quasilinear $(p,q)$-sawyerable operator $T$ is bounded from $X(\mathcal{R},\mu)$ to $Y^{\langle p\rangle}(\mathcal{S},\nu)$.
@@
Let $B$ be a linear operator defined at least on characteristic functions of $\mu$-measurable subsets of $\mathcal{R}$ of finite measure and taking values in $\mathfrak{M}_{0}(S,\nu)$. Let $G$ be a $(\lambda p,\lambda q)$-sawyerable quasilinear operator. Assume that $R_{p,q}\colon\bar{X}(0,\mu(\mathcal{R}))\to\bar{Y}(0,\nu(\mathcal{S}))$, and that there is a constant $C>0$ such that for every $\mu$-measurable set $E\subseteq\mathcal{R}$ of finite measure:
@@
The usefulness of Theorems 1.3 and 1.4 can be fully understood in detail by the applications discussed at the end of Section 3. Their important application involving Lorentz spaces, which generalizes Theorem A, follows. In fact, it is a corollary of a more general theorem, namely Theorem 3.7, which establishes a general boundedness result for the Riesz potential $I_{\alpha}^{\mu}$, defined by (1.27), under suitable assumptions on a pair of Radon measures $\mu,\nu$. It relies on our principle to connect a bad operator, $I_{\alpha}^{\mu}$, to a good one, for which bounds are known. Here the good operator is a suitable dyadic maximal function associated to the measure $\mu$ for which bounds were established by E. Sawyer in [Sawyer1] (see the discussion before Lemma 3.5 for more details).
Theorem 1.6.
(1.26) $\sup_{x\in\mathbb{R}^{n},r>0}\frac{\mu(B(x,r))}{r^{d}}<\infty$
and
$\sup_{\mu(Q)>0}\frac{\nu(Q)}{\mu(Q)^{1-\frac{\alpha p}{d}}}<\infty.$
When $\mu$ is the Lebesgue measure on $\mathbb{R}^{n}$ (and so $d=n$), this recovers the Korobkov–Kristensen trace inequality (1.5) by a different argument. Consequently, we obtain the assertion of Theorem A (hence also Theorem B) as a special case of our results.
@@ 2. Preliminaries
(2.3) $\int_{E}|f(x)|\,\mathrm{d}{\mu(x)}\leq\int_{0}^{\infty}(f\chi_{E})^{*}_{\mu}(t)\,\mathrm{d}{t}\quad\text{for every $f\in\mathfrak{M}(\mathcal{R},\mu)$}$
@@
When $\mu(\mathcal{R})<\infty$, $X^{\langle\alpha\rangle}(\mathcal{R},\mu)$ is always a rearrangement-invariant function space. When $\mu(\mathcal{R})=\infty$, $X^{\langle\alpha\rangle}(\mathcal{R},\mu)$ is a rearrangement-invariant function space unless it is trivial (i.e., it contains only the zero function, and so it does not satisfy (P4)), which may happen. For a detailed study of the spaces $X^{\langle\alpha\rangle}(\mathcal{R},\mu)$, see [Tur:23] (recall also (1.24)).
@@ 3. Sawyerability and properties of the governing operator
Let $(\mathcal{R},\mu)$ and $(\mathcal{S},\nu)$ be nonatomic $\sigma$-finite measure spaces. Assume that $p,q$ satisfy (1.12), and let $r$ be defined by (1.15). Let $T$ be a quasilinear operator defined on $(L^{p}+L^{q,\infty})(R,\mu)$ and taking values in $\mathfrak{M}_{0}(S,\nu)$. Then the following three statements are equivalent.
@@
First, assume that (i) is true, i.e., $T$ is bounded from $L^{p}(\mathcal{R},\mu)$ and $L^{q,\infty}(\mathcal{R},\mu)$ to $L^{p}(\mathcal{S},\nu)$ and $L^{\infty}(\mathcal{S},\nu)$, respectively. Since $T$ is quasilinear, it follows (see [BK:91, Proposition 3.1.15] and [BS, Chapter 5, Theorem 1.11]) that (3.2) is true with a constant $C>0$ depending only on $\|T\|_{L^{p}\to L^{p}}$, $\|T\|_{L^{q,\infty}\to L^{\infty}}$, and the constant $k$ from the definition of the quasilinearity. In other words, (i) implies (iii).
@@
We now turn our attention to specializing Theorems 1.3 and 1.4. We start with customizing Theorem 1.4 to the generalized potential $I_{\alpha}^{\mu}$ defined by (1.27), which will serve as the “bad operator” $B$ in (1.19). For this “bad operator”, the “good operator” $G$ is a suitable fractional maximal function (or rather, a suitable family of fractional maximal functions).
@@
Here $\mathcal{H}^{0}$ is the counting measure on $\mathbb{R}^{n}$. Furthermore, it is not hard to see that our assumptions on the measures $\mu$ and $\nu$ imply that
@@
Then every quasilinear $(p,q)$-sawyerable operator $T$ is bounded from $L^{r_{1},s_{1}}(\mathcal{R},\mu)$ to $L^{r_{2},s_{2}}(\mathcal{S},\nu)$.
@@
Finally, assume that $\nu(\mathcal{S})<\infty$ and either (3.11) or (3.12) is valid. We have already proved that every quasilinear $(p,q)$-sawyerable operator $T$ is bounded from $L^{r_{1},s_{1}}(\mathcal{R},\mu)$ to $L^{r_{2},s_{2}}(\mathcal{S},\nu)$. Since $L^{r_{2},s_{2}}(\mathcal{S},\nu)\subseteq L^{r,s}(\mathcal{S},\nu)$ for every $1\leq r<r_{2}$ and $s\in[1,\infty]$ provided that $\nu(\mathcal{S})<\infty$ (e.g., [BS, p. 217]), we immediately obtain the fact that $T\colon L^{r_{1},s_{1}}(\mathcal{R},\mu)\to L^{r,s}(\mathcal{S},\nu)$ is also bounded. ∎
@@
Then every linear operator $B$, defined at least on characteristic functions of $\mu$-measurable subsets of $\mathcal{R}$ of finite measure and taking values in $\mathfrak{M}_{0}(S,\nu)$, satisfying (1.19) with some $(\lambda p,\lambda q)$-sawyerable quasilinear operator $G$, is bounded from $L^{r_{1},1}(\mathcal{R},\mu)$ to $L^{r_{2},s_{2}}(\mathcal{S},\nu)$.
@@ 4. Proofs of Main Results
We start by showing that (iii) implies (ii). Let $T$ be a quasilinear $(p,q)$-sawyerable operator. By Proposition 3.1, one has
@@
As (ii) clearly implies (i), we only need to prove that (i) implies (iii). Fix $g=g\chi_{(0,\mu(\mathcal{R}))}\in\bar{X}(0,\mu(\mathcal{R}))$. Replacing $g$ with $g\chi_{(0,N)}\chi_{(0,\mu(\mathcal{R}))}$ for an appropriate $N\in(0,\infty)$ if necessary, we may assume that $|\operatorname{supp}g|<\infty$. Since $(R,\mu)$ is nonatomic, there is a function $h\in X(R,\mu)\subseteq(L^{p}+L^{q,\infty})(R,\mu)$ such that (see [BS, Chapter 2, Corollary 7.8])
@@
We first note that, by Lemma 4.1, the operator $G_{\frac{1}{\lambda}}$ is $(p,q)$-sawyerable. Furthermore, it is easy to see that the quasilinearilty of $G$ implies that $G_{\frac{1}{\lambda}}$ is also quasilinear. Consequently, since $R_{p,q}\colon\bar{X}(0,\mu(\mathcal{R}))\to\bar{Y}(0,\nu(\mathcal{S}))$, Theorem 1.3 implies that
@@
for (4.26) to possibly hold. A characterization of the validity of (4.26) under the restriction (4.27) is provided by [Pic:13, Theorem 10.3.12, (i)–(ii)]. It read as: either $s_{1}>p$,

### v2

@@ Abstract.
In this paper we develop a general theoretical tool for the establishment of the boundedness of notoriously difficult operators (such as potentials) on certain specific types of rearrangement-invariant function spaces from analogous properties of operators that are easier to handle (such as fractional maximal operators). A principal example of the new results one obtains by our analysis is the following inequality, which generalizes a result of Korobkov and Kristensen (who had treated the case $\mu=\mathcal{L}^{n}$, the Lebesgue measure on $\mathbb{R}^{n}$): There exists a constant $C>0$ such that
$\int_{\mathbb{R}^{n}}|I_{\alpha}^{\mu}f|^{p}\,\mathrm{d}{\nu}\leq C\|f\|_{L^{p,1}(\mathbb{R}^{n},\mu)}^{p}$
for all $f$ in the Lorentz space $L^{p,1}(\mathbb{R}^{n},\mu)$, where $\mu,\nu$ are Radon measures such that
$\sup_{x\in\mathbb{R}^{n},r>0}\frac{\mu(B(x,r))}{r^{d}}<\infty\quad\text{and}\quad\sup_{\mu(Q)>0}\frac{\nu(Q)}{\quad\mu(Q)^{1-\frac{\alpha p}{d}}}<\infty,$
and $I_{\alpha}^{\mu}$ is the Riesz potential defined with respect to $\mu$ of order $\alpha\in(0,d)$. More broadly, we obtain inequalities in this spirit in the context of rearrangement-invariant spaces through a result of independent interest, an extension of an interpolation theorem of Calderón where the target space in one endpoint is a space of bounded functions.
@@
Trace inequality, Hausdorff content, interpolation of operators, Calderón theorem, rearrangement-invariant spaces, Lorentz spaces, potentials, fractional maximal function
@@ 1. Introduction and statement of main results
using the Hahn-Banach theorem one obtains that
@@
for every $f\in L^{p,1}(\mathbb{R}^{n})$. One can replace the $\approx$ with equality in this statement about Banach space duality by utilizing a norm on $L^{1}(\mathcal{H}^{n-\alpha p}_{\infty})$ in place of the above quasi-norm, for example with a Choquet integral involving the corresponding dyadic Hausdorff content. Now, while Theorem A asserts that the left-hand side of (1.6) is bounded from above by a constant multiple of $\|f\|_{L^{p,1}(\mathbb{R}^{n})}^{p}$, [Kristensen-Korobkov, Theorem 1.2] asserts the same for the right-hand side of (1.6). Therefore, the formulations are indeed equivalent.
@@
The starting point of this paper is an observation concerning the connection of the result of Korobkov and Kristensen and a classical result of E. Sawyer [Sawyer1, Sawyer2], combined with the same duality principle as in Remark 1.1. To this end, we recall that in his papers on one and two weight estimates, Sawyer proved (see also [Ada:98, pp. 28–29]) the following: For $\beta\in(0,n)$ and $1<q<\frac{n}{\beta}$ there exists a constant $C_{7}=C_{7}(\beta,q,n)>0$ such that
@@
For $\beta=\alpha$ and $q=p$, the inequality (1.4) shows that the estimate for the fractional maximal function in (1.8) is better than that for the Riesz potential of the same order in (1.1), while, as discussed in the preceding, there is no hope to control the integral of the Riesz potential with respect to the Hausdorff content of this order.
@@
Thus, for any $\alpha\in(0,n)$ and $1<p<\frac{n}{\alpha}$, one may choose $\beta\in(\alpha,n)$ such that $q=p\frac{\alpha}{\beta}>1$. Then, for $f=\chi_{E}$ and $\theta=\alpha/\beta$, (1.9) yields
@@
This approach seems to have gone unnoticed until now, despite the great interest in trace inequalities, and is a key idea for the new results we establish in this paper. In particular, in this work we develop a framework for this principle, when a bound for a good operator like the fractional maximal operator can easily be translated into a bound for a (comparatively) bad operator like the Riesz potential. To demonstrate the new theory we develop with a concrete example, we record here the following generalization of Korobkov and Kristensen’s result.
Theorem 1.2.
(1.11) $\sup_{x\in\mathbb{R}^{n},r>0}\frac{\mu(B(x,r))}{r^{d}}<\infty$
and
(1.12) $\sup_{\mu(Q)>0}\frac{\nu(Q)}{\mu(Q)^{1-\frac{\alpha p}{d}}}<\infty.$
When $\mu$ is the Lebesgue measure on $\mathbb{R}^{n}$ (and so $d=n$), this recovers the Korobkov–Kristensen trace inequality (1.5) by a different argument. In particular, our Theorem 1.2 and the argument presented in Remark 1.1 yield the assertion of Theorem A (hence also Theorem B) as a special case of our results. In this case, and in general with (1.11) as given, the assumption (1.12) is necessary, which can be seen by taking $f=\chi_{Q}$. However, in general the combination of (1.11) and (1.12) need not be necessary and may be relaxed to the imposition of the testing condition of E. Sawyer in [Sawyer1] which our work builds upon.
Theorem 1.2 provides a specific example of the new results which follow from our work, though our results hold in the broader context of rearrangement-invariant function spaces (see Theorem 3.7). This preempts the question of bounds for bad operators with that of necessary and sufficient conditions for bounds for good operators on these spaces. This general context of rearrangement-invariant function spaces provides a unifying theory for function spaces such as Lebesgue spaces, Lorentz spaces, or Orlicz spaces, to name a few.
@@
Recall that the corresponding restricted weak-type estimates for a quasi-linear operator $T$ read as:
$T\colon L^{p,1}(\mathbb{R}^{n},dx)\to L^{p,\infty}(\mathbb{R}^{n},\nu),$
$T\colon L^{\frac{n}{\alpha},1}(\mathbb{R}^{n},dx)\to L^{\infty}(\mathbb{R}^{n},\nu).$
When $T=M_{\alpha}$, the conclusion of the classical Calderón theorem is not optimal, because it does not fully exploit the endpoint boundedness properties (1.14) and (1.15), which are better than of restricted weak-type. As we will see, this results in significant differences in the theory we develop from the classical one. Several other types of various nonstandard versions of Calderón’s theorem can be found in literature, see e.g. [Bae:22, GP-Indiana:09, Mal:12], but none of the known ones can be used for our purposes.
@@
We say that a quasi-linear operator $T$ defined on $(L^{p}+L^{q,\infty})(\mathcal{R},\mu)$ and taking values in $\mathfrak{M}_{0}(\mathcal{S},\nu)$ is $(p,q)$-sawyerable if $T$ satisfies:
@@
Recall that an operator $T$ defined on a linear space $X\subseteq\mathfrak{M}_{0}(R,\mu)$ and taking values in $\mathfrak{M}_{0}(S,\nu)$ is quasi-linear if there is a constant $k\geq 1$ such that
@@
for every $f,g\in X$ and every scalar $\alpha$. We say that $T$ is sublinear if it is quasi-linear with $k=1$.
@@
    Every quasi-linear $(p,q)$-sawyerable operator $T$ is bounded from $X(\mathcal{R},\mu)$ to $Y^{\langle p\rangle}(\mathcal{S},\nu)$.
@@
Let $B$ be a linear operator defined at least on characteristic functions of $\mu$-measurable subsets of $\mathcal{R}$ of finite measure and taking values in $\mathfrak{M}_{0}(S,\nu)$. Let $G$ be a $(\lambda p,\lambda q)$-sawyerable quasi-linear operator. Assume that $R_{p,q}\colon\bar{X}(0,\mu(\mathcal{R}))\to\bar{Y}(0,\nu(\mathcal{S}))$, and that there is a constant $C>0$ such that for every $\mu$-measurable set $E\subseteq\mathcal{R}$ of finite measure:
@@
The strength and generality of Theorems 1.4 and 1.5 can be fully understood in detail by the applications discussed at the end of Section 3. Their important application involving Lorentz spaces, which generalizes Theorem A, is our Theorem 1.2 stated above. In fact, it is a corollary of a more general theorem, namely Theorem 3.7, which establishes a general boundedness result for the Riesz potential $I_{\alpha}^{\mu}$, defined by (1.13), under suitable assumptions on a pair of Radon measures $\mu,\nu$. It relies on our principle to connect a bad operator, $I_{\alpha}^{\mu}$, to a good one, for which bounds are known. Here the good operator is a suitable dyadic maximal function associated to the measure $\mu$ for which bounds were established by E. Sawyer in [Sawyer1] (see the discussion before Lemma 3.5 for more details).
@@ 2. Preliminaries
(2.3) $\int_{E}|f(x)|\,\mathrm{d}{\mu(x)}\leq\int_{0}^{\mu(E)}f^{*}_{\mu}(t)\,\mathrm{d}{t}\quad\text{for every $f\in\mathfrak{M}(\mathcal{R},\mu)$}$
@@
When $\mu(\mathcal{R})<\infty$, $X^{\langle\alpha\rangle}(\mathcal{R},\mu)$ is always a rearrangement-invariant function space. When $\mu(\mathcal{R})=\infty$, $X^{\langle\alpha\rangle}(\mathcal{R},\mu)$ is a rearrangement-invariant function space unless it is trivial (i.e., it contains only the zero function, and so it does not satisfy (P4)), which may happen. For a detailed study of the spaces $X^{\langle\alpha\rangle}(\mathcal{R},\mu)$, see [Tur:23] (recall also (1.28)). Let us recall that these spaces play a decisive role for Sobolev embeddings into spaces with slowly decaying Frostman measures, as was recently pointed out in [Cia:20, Theorem 5.1].
@@ 3. Sawyerability and properties of the governing operator
Let $(\mathcal{R},\mu)$ and $(\mathcal{S},\nu)$ be nonatomic $\sigma$-finite measure spaces. Assume that $p,q$ satisfy (1.16), and let $r$ be defined by (1.19). Let $T$ be a quasi-linear operator defined on $(L^{p}+L^{q,\infty})(R,\mu)$ and taking values in $\mathfrak{M}_{0}(S,\nu)$. Then the following three statements are equivalent.
@@
First, assume that (i) is true, i.e., $T$ is bounded from $L^{p}(\mathcal{R},\mu)$ and $L^{q,\infty}(\mathcal{R},\mu)$ to $L^{p}(\mathcal{S},\nu)$ and $L^{\infty}(\mathcal{S},\nu)$, respectively. Since $T$ is quasi-linear, it follows (see [BK:91, Proposition 3.1.15] and [BS, Chapter 5, Theorem 1.11]) that (3.2) is true with a constant $C>0$ depending only on $\|T\|_{L^{p}\to L^{p}}$, $\|T\|_{L^{q,\infty}\to L^{\infty}}$, and the constant $k$ from the definition of the quasi-linearity. In other words, (i) implies (iii).
@@
We now turn our attention to specializing Theorems 1.4 and 1.5. We start with customizing Theorem 1.5 to the generalized potential $I_{\alpha}^{\mu}$ defined by (1.13), which will serve as the “bad operator” $B$ in (1.23). For this “bad operator”, the “good operator” $G$ is a suitable fractional maximal function (or rather, a suitable family of fractional maximal functions).
@@
Here $\mathcal{H}^{0}$ is the counting measure on $\mathbb{R}^{n}$ and $L^{\infty}(\mathbb{R}^{n},\mathcal{H}^{0})$ is the space of everywhere bounded functions. Furthermore, it is not hard to see that our assumptions on the measures $\mu$ and $\nu$ imply that
@@
Then every quasi-linear $(p,q)$-sawyerable operator $T$ is bounded from $L^{r_{1},s_{1}}(\mathcal{R},\mu)$ to $L^{r_{2},s_{2}}(\mathcal{S},\nu)$.
@@
Finally, assume that $\nu(\mathcal{S})<\infty$ and either (3.11) or (3.12) is valid. We have already proved that every quasi-linear $(p,q)$-sawyerable operator $T$ is bounded from $L^{r_{1},s_{1}}(\mathcal{R},\mu)$ to $L^{r_{2},s_{2}}(\mathcal{S},\nu)$. Since $L^{r_{2},s_{2}}(\mathcal{S},\nu)\subseteq L^{r,s}(\mathcal{S},\nu)$ for every $1\leq r<r_{2}$ and $s\in[1,\infty]$ provided that $\nu(\mathcal{S})<\infty$ (e.g., [BS, p. 217]), we immediately obtain the fact that $T\colon L^{r_{1},s_{1}}(\mathcal{R},\mu)\to L^{r,s}(\mathcal{S},\nu)$ is also bounded. ∎
@@
Then every linear operator $B$, defined at least on characteristic functions of $\mu$-measurable subsets of $\mathcal{R}$ of finite measure and taking values in $\mathfrak{M}_{0}(S,\nu)$, satisfying (1.23) with some $(\lambda p,\lambda q)$-sawyerable quasi-linear operator $G$, is bounded from $L^{r_{1},1}(\mathcal{R},\mu)$ to $L^{r_{2},s_{2}}(\mathcal{S},\nu)$.
@@ 4. Proofs of Main Results
We start by showing that (iii) implies (ii). Let $T$ be a quasi-linear $(p,q)$-sawyerable operator. By Proposition 3.1, one has
@@
As (ii) clearly implies (i), we only need to prove that (i) implies (iii). Fix $g=g\chi_{(0,\mu(\mathcal{R}))}\in\bar{X}(0,\mu(\mathcal{R}))$. Replacing $g$ with $g\chi_{(0,N)}\chi_{(0,\mu(\mathcal{R}))}$ for an appropriate $N\in(0,\infty)$ if necessary, we may assume that $|\operatorname{supp}g|<\infty$, i.e. the support of $g$ has finite measure. Since $(R,\mu)$ is nonatomic, there is a function $h\in X(R,\mu)\subseteq(L^{p}+L^{q,\infty})(R,\mu)$ such that (see [BS, Chapter 2, Corollary 7.8])
@@
We first note that, by Lemma 4.1, the operator $G_{\frac{1}{\lambda}}$ is $(p,q)$-sawyerable. Furthermore, it is easy to see that the quasi-linearilty of $G$ implies that $G_{\frac{1}{\lambda}}$ is also quasi-linear. Consequently, since $R_{p,q}\colon\bar{X}(0,\mu(\mathcal{R}))\to\bar{Y}(0,\nu(\mathcal{S}))$, Theorem 1.4 implies that
@@
for (4.26) to possibly hold. A characterization of the validity of (4.26) under the restriction (4.27) is provided by [Pic:13, Theorem 10.3.12, (i)–(ii)]. It reads as: either $s_{1}>p$,